7.42 What is the de Broglie wavelength (in nm) associ- iated with a 2.5-g Ping-Pong ball traveling 35 mph?
Answer #1

Given:

Mass of the Ping-Pong ball = 2.5 g

Speed of the Ping-Pong ball = 35 mph

STEP 1: Calculate the mass in SI units

m=2.5 g

m=2.5 \times 10^{-3} kg   (\because 1g=10^{-3}kg)

STEP 2: Calculate the speed in SI units

v=35 mph

V = 35 X 0.447m/s

v=15.645 m/s   (\because 1mph = 0.447 m/s)

STEP 3: Use the de- Broglie equation to calculate the wavelength of the ball

\lambda =\frac{h}{mv}   where h is the Planck's Constant

\lambda =\frac{6.626\times 10^{-34}Js}{(2.5\times 10^{-3}kg)\times (15.645 m/s)}

\lambda =1.69 \times 10^{-32} m

STEP 4: Convert meters into nanometers

\because 1m=10^{9} nm

\therefore \lambda =1.69 \times 10^{-32} \times 10^{9} nm

\lambda =1.69 \times 10^{-23} nm

Therefore, the wavelength associated with the Ping-Pong ball is 1.69 \times 10^{-23} nm